# Setting the Course, but Forgetting to Steer: Analyzing Compliance with GDPR’s Right of Access to Data by Instagram, TikTok, and YouTube

### v1

@@ 3.3 Information regarding user’s content
date, and time information
@@
Table 3: Details of user created content shared by platforms.
@@
While all three platforms provide a copy of the media and text details, Instagram shares additional information including software used to upload the content, device ID, metadata about the camera, etc. (Table 3).
3.4 Information regarding personal details
@@
While the two facets discussed above focus more on the content that users consume, engage with, or create, the information regarding personal details is more privacy-sensitive. These details contain basic personally identifiable information (PII), e.g., account details (which may include one’s name, phone number, e-mail address, date of birth, profile picture, etc.). However, apart from these basic personal details, most of the DDPs also include information about the different devices through which people visit the platforms. TikTok and Instagram directly store these details as part of the login information, whereas YouTube does not offer this data in its DDP. For YouTube, this data can be accessed only if someone requests their entire Google DDP. [^2] One justification may be that one can not sign in to their YouTube account unless they sign in to their Google account on the same device.
@@
Login history: In light of the prior descriptions, login history contains the list of login activities made by the data subject. The login history details are significantly different in TikTok and Instagram DDPs. TikTok provides the timestamp, IP, device model, operating system, network type, and carrier provider in its DDP. On the other hand, Instagram’s login history data is very detailed. Along with providing the above details, Instagram provides (and therefore collects) cookie information, language code, Instagram app version, display properties of the device, hardware identifier, and some internal identifiers. Such differences in data collection and sharing practices point to two important possibilities: (a) TikTok might be collecting some of these details and not sharing them with its end-users and, therefore, is not being transparent; (b) Instagram might be collecting more information than necessary, thus becoming more intrusive to privacy. Diving deep into these possibilities is beyond the scope of this study, and we would like to explore these directions in future work.
3.5 Information regarding advertisements
@@
Advertisement on social media platforms is a prevalent phenomenon, as users are constantly served with ads when they scroll through Instagram, TikTok, or YouTube. However, based on our observations, there is no standardized way of including ads that users have seen in their requested DDP.
Ads viewed: As mentioned in Section 3.2, Instagram maintains a separate list of ads viewed by users. However, the data provided by Instagram only mentions the author/advertiser name and timestamp. TikTok, on the other hand, does not demarcate any advertised content that a user has seen in their timeline. In contrast, YouTube distinguishes all the ads that a user has encountered from the organic content. Also, YouTube provides a link to the ad video along with its title and when it was viewed. However, neither of the three platforms provides information on ad targeting parameters. From the ads viewed details, a user can understand these are the ads they got to watch on their timeline, but why those particular ads were served to them remains an enigma. While this information does appear when an advertisement or paid content appears on the three platforms, it is not included in the shared DDPs.
@@
Ad targeting data sources: The Instagram DDP contains a list of advertisers who have used data about one’s online activity or profile for targeted advertising purposes. This list also conveys information about how these advertisers obtained users’ data. In specific, as per this information, there are three ways in which the advertiser might target a person: (a) custom audience – where the advertiser targets the data subject through a custom audience with a list of customer data (e.g., email addresses, phone numbers, or other identifiers) which the advertiser might have obtained from other third party sources; (b) remarketing custom audience – where the advertiser targets the data subject because (s)he might have interacted with some of the advertiser’s content, visited their website, etc.; (c)  in-person store visit – where advertiser can use location data or check in data because the data subject might have visited one of their physical stores. The DDPs of YouTube and TikTok do not provide any such information on advertisers and how they gather data about a particular user.
Off-platform activities: Another important source of data through which these platforms collect enormous amounts of digital behavioral data is off-platform activities. This is a list of all the activities in external platforms (outside the source platform) that these platforms have tracked or linked to a user. Note that this might be an artifact of the said external platforms’ installed Meta/Google/TikTok pixels. A user might have used a Meta or Google login to sign up, other tools embedded in the software development kit of the external platform, or any data-sharing agreement between the external company and any of the platforms we investigate.
@@
For instance, if a user visits the website or application that also installs one of the tracking techniques of any of the short-format platforms, then the said website can share the user’s activities, such as page view, purchase, search, logins, etc., along with other activities, such as adding a product to the cart, view contents, install app, launch app, etc. Instagram and TikTok DDPs contain off-platform activities of users, whereas YouTube does not have such data. Both Instagram and TikTok share the external platform name, event or activity type, and the timestamp in their DDPs.
@@
The DDPs from the three platforms also provide a wide variety of other details. For example, they may contain information on link history, survey, application setting, shopping, etc. For more details on this, please refer to the summary of all the discussions above, listed in Table 5 (see Appendix A). All three platforms miss one of the important goals set by GDPR Article 15(1)(a), which is to include the purposes of the data processing and collection. None of the DDPs include a purpose statement explaining why certain data is collected, nor do they provide a README or manual to help users understand the significance of the data being collected.
Important takeaways
[…]
  The three platforms implement Article 15 very differently. The Instagram and TikTok DDPs contain a significantly larger number of data types than the YouTube DDP. Instagram shares (and thus collects) more data.
@@
4 RQ2(a): How comprehensible are DDPs?
Here we investigate whether the current implementations of platforms adhere to the requirements Chapter 3 of the GDPR outlines a number of rights of data subjects (i.e., end users), including the right of access (Article 15) [12]. Before outlining these rights, the GDPR provides a list of requirements for how information should be communicated to end users. Specifically, the first paragraph of Article 12 of GDPR reads, “The controller shall take appropriate measures to provide any information referred to in Articles 13 and 14 and any communication under Article 15 to 22 and 34 relating to processing to the data subject in a concise, transparent, intelligible and easily accessible form, using clear and plain language $\dots$ [12].
@@ 4 RQ2(a): How comprehensible are DDPs?
As the GDPR does not provide more succinct definitions of these requirements, we adopt two approaches to interpret these terms. First, following trends in empirical contract research to employ user surveys for interpretation of legal terms [5], we conduct a large-scale user survey among participants from different European countries to understand their interpretations of some of the requirements. Second, we apply the interpretations of the requirements by the European Data Protection Board (EDPB) [14]. Finally, we evaluate the current implementations, i.e., the DDPs obtained from Instagram, TikTok, and YouTube, against the interpretations by users and the EDPB.
@@
Participant recruitment: We recruited participants to conduct our survey from the crowd-sourcing platform Prolific[27]. On Prolific, we hired 100 participants each from Germany, France, Spain, and Italy (400 in total) who had a high approval rate of at least 98% to take part in our survey. Our motivation for a general set of respondents stems from the fact that the GDPR entitles general residents in the European Union(EU) with the right to data access, and requirements are intended to benefit these residents. Though, eventually, our plan is to extend the study to all countries within the EU, we chose to start with the countries that have the highest GDP [32]. We also selected the standard gender breakdown while setting up the surveys. In total, we recruited 228 males (57.0%), 161 females (40.3%), 9 self-reported as “other” (2.3%), and 2 participants who preferred not to disclose their gender (0.5%) (See Table 6 in Appendix C for details).
Prior to entering our survey, we presented an online consent form for expressing explicit consent from the participants. Upon completion of the survey, we compensated each participant at a rate of £9 per hour, which is recommended by Prolific to be a good and ethical rate of remuneration [26]. Our survey comprised three components: (a) Awareness of participants about the rights, (b) Interpretation of the requirements by the participants and (c) adherence evaluation. Next, we elaborate abut the survey setup and important observations for each of the components.
@@ 4.2 Interpretation of the requirements
Survey setup: To understand the interpretation of requirements by common end users, in the second part of our survey, we showed the participants the paragraph of Article 12 (see above), which mentions the different requirements of DDPs. We asked our participants about their interpretation or expectation of the shared personal data in the DDPs to follow any of these requirements. Next, we elaborate on the different interpretations of these requirements from the responses of our participants while contextualizing them further with the guidelines of the EDPB [14], which represents the data protection authorities of all EU member states.
Conciseness Interpretations by the survey participants: We perform thematic analysis of the responses provided by the 400 participants using BERTopic [16]. The important themes that resonated across the responses spanning all countries are – prioritize important & relevant data, summarize important observations, and present data in a structured way that is easier to navigate.
However, a fourth theme emerged in the interpretations shared by participants from Italy, France, and Spain, but not by those from Germany. Specifically, these participants touched upon the concept of data minimization. To this end, some participants mentioned the documents should be shorter, while others mentioned avoiding an overwhelming amount of text. In short, while participants from Germany only prioritize the informativeness of the content, other participants prioritize the length of the content as well.
Interpretations by the EDPB: In this context, the European Data Protection Board interprets concise presentation to be done efficiently and succinctly to avoid information fatigue [14]. While the above interpretation falls within the interpretations mentioned by our survey participants, the EDPB also asserts that the presentation should differentiate privacy-sensitive information from non-privacy-sensitive information.
@@
Transparency Interpretation by the survey participants: Unlike the case of conciseness, there are no major differences in how participants from all countries interpret transparency. The interpretation of transparency can be divided into two major aspects. First, participants call for unambiguous, full disclosure of all the data that has been collected about them in the way they are collected. Second, and most importantly, some participants also expect the DDPs to mention how the data is processed and for what purpose the data is collected. To this extent, the response from one of the participants captures all the notions beautifully: “It (Transparency) means providing clear honest and complete information about how the data is being used. There should be no hidden details or surprises for the individual. People should easily understand what data is being collected why its being collected who will access it and how long it will be kept.” At this point, we would also like to note that as per the participants’ interpretation, conciseness and transparency are two competing requirements, i.e., to satisfy one, the other possibly has to be violated.
Interpretation by the EDPB: While the EDPB draws most of its interpretations from Recital 39 of the GDPR, it interprets transparency as allowing data subjects to determine the scope and consequences of the concerned data processing [14]. To this end, the EDPB calls for data controllers (i.e., platforms) to separately spell out the most important consequences and risks that the data processing may entail.
@@
Intelligibility Interpretation by the survey participants: Among all the requirements provided by the GDPR, the most subjective one is that of intelligibility. However, almost all the participants explicitly interpret it to be easy to understand and link it closely with the language being straightforward and without any technical and/or legal jargon. Participants from all countries also mentioned that intelligible data should have a clear explanation of how and why a certain piece of information is collected. The participants also noted some interesting expectations in this regard. For instance, participants from Germany mentioned that platforms shall share their DDPs in one’s language of comfort (native language or other language of choice). Moreover, participants from Italy stated that platforms should share their DDPs in the form of elegant visualization.
Interpretation by the EDPB: The EDPB describes intelligibility to be a representation that should be understood by an average member of the intended audience. In addition, they also mention that it is closely linked to the requirement to use clear and plain language.
@@
Clear and plain language Interpretation by the survey participants: Understandably, most of the participants interpret plain and simple language to be words that are easily understandable and unambiguous. Some participants elaborate on the nuance by stating that legal and technical terms should be avoided or at least explained in a simpler way. To this end, one clear observation from our survey responses is that participants do not distinguish between intelligibility and language.
Interpretation by the EDPB: Along the lines of the interpretations mentioned above, the EDPB interprets this requirement as being definitive and should not be phrased in ambivalent terms or leave room for different interpretations [14].
@@
Accessibility Interpretation by the survey participants: Most of our participants interpret accessibility as being easily available, i.e., it is easy to find the place to request the data and to access the data. However, there are certain interesting and valid interpretations which are worth mentioning. Some participants mentioned that the data needs to be available in a format so that it can be opened without any sophisticated software or tools. At the same time, the timeliness at which the data is provided upon request is another interesting aspect touched upon by some of our participants. Another interesting interpretation was the data should be available to everyone, including people with visual impairments by using accessible design features like screen readers or alternative text making sure the users can access data without any barriers.
Interpretation by the EDPB: The EDPB’s interpretation of accessibility also follows from the above popular responses from the survey participants, wherein they say that it should be immediately apparent to end-users as to where and how the DDP can be accessed [14].
Our observations on accessibility: Despite agreeing with all of the above interpretations, we observe a few differences across the platforms because accessibility is the most objective property for evaluation. In terms of ease of request, from the content page of a user’s account, one needs to click 6, 6, and 10 times on TikTok, YouTube, and Instagram, respectively, to be able to request the data in the human-readable format. Notice that, while on Instagram, a user needs to click more, the platform provides many options to the end-user, such as the time duration that their DDP should contain. Moreover, based on our anecdotal observations, the turnaround time for TikTok upon request was instantaneous during this research study. For Instagram and YouTube, it often takes 10 to 15 minutes for the DDP to become available. However, all the turnaround times are within the GDPR-prescribed maximum duration of one month. Further, one can download one’s data on TikTok and Instagram from both their applications and web versions, but YouTube shares the DDP with its end users by sending an e-mail with a URL to the DDP. All these platforms provide DDPs in a compressed format, which makes it harder for the end-users to open them on a smartphone. At the same time, an end user needs to have a browser or text reader to open the HTML (YouTube and Instagram) and TXT (TikTok) DDPs. In summary, while the data request and download are relatively accessible in the current setting, reading the data is still complicated for the average end user. At this point, we would also like to note that we are not aware of any special provisions that platforms may have for people with visual impairments to access the data.

### v2

@@ 3.3 User’s content
date and time information
@@
TABLE IV: User created content shared by platforms.
@@
While all three platforms provide a copy of the media and text details, Instagram shares additional information including software used to upload the content, device ID, metadata about the camera, etc. (Table IV).
3.4 Personal details
@@
Distinct from data on content engagement, information regarding personal details is inherently more sensitive. These details contain basic personally identifiable information (PII), e.g., account details, one’s name, phone number, e-mail address, date of birth, profile picture, etc. However, apart from these basic details, most of the DDPs also include a lot of other sensitive information about the users.
@@
Login history: Login history contains the list of login activities made by the data subject. While TikTok and Instagram provide these details in their DDPs, YouTube does not provide them in its DDP. For YouTube, these details are usually found in the Google DDP. The login history details are significantly different in TikTok and Instagram DDPs. While TikTok provides the timestamp, IP, device model, operating system, network type, and carrier provider in its DDP, Instagram’s login history data is significantly more detailed. Along with the above details, Instagram provides (and therefore collects) cookie information, language code, Instagram app version, display properties of the device, hardware identifier, and some internal identifiers.
Such differences in the amount of PII shared point to two important possibilities: (a) TikTok might be collecting some of these details and not sharing them with its end-users and, therefore, is not being transparent; or (b) While such detail might be argued as relevant for security purposes, the extent of Instagram’s collection, particularly regarding persistent identifiers and granular device specifics, warrants scrutiny under data minimization principles (Article 5(1)(c)). Investigating these is beyond the scope of this study, and we would like to explore these directions in future work.
3.5 Advertisements
@@
Users are constantly served with ads when they scroll through TikTok, Instagram, and YouTube. However, based on our observations, there is no standardized way of including ads that users have seen in their requested DDP.
Ads viewed: While YouTube and Instagram demarcate ads viewed by users, TikTok does not distinguish them in its DDP. Among the two platforms that provide ads data, Instagram maintains it in a separate list, whereas YouTube labels ads in its watch history file. YouTube provides ads’ title, URL, advertiser details, and timestamp; in contrast, Instagram only shares the advertiser name and timestamp.
However, neither of the three platforms provides information on ad targeting parameters. From the ads viewed details, a user can understand these are the ads they viewed on their timeline, but why those particular ads were served to them remains an enigma. This lack of transparency regarding ad targeting could be seen as not fully aligning with the spirit of Article 15(1)(h) concerning meaningful information about automated decision-making, including profiling.
@@
Ad targeting data sources: The Instagram DDP contains a list of advertisers who have used data about one’s online activity or profile for targeted advertising purposes, specifically to reach them as an identified individual (as opposed to targeting based on general demographics or interests). This list also conveys information about how these advertisers obtained users’ data. In specific, as per this information, there are three ways in which the advertiser targets a user: (a) custom audience – where the advertiser targets the user through a custom audience with a list of customer data (e.g., email addresses, phone numbers, or other identifiers) which the advertiser might have obtained from other third party sources; (b) remarketing custom audience – where the advertiser targets the user because (s)he might have interacted with some of the advertiser’s content, visited their website, etc.; (c)  in-person store visit – where advertiser can use location data or check in data because the data subject might have visited one of their physical stores.
YouTube and TikTok DDPs do not provide any such information on advertisers and how they gather data about a particular user. Given that ad targeting often involves processing data from various sources, this omission may be a potential violation of Article 15(1)(g) which mandates disclosing such sources and the gathered data.
@@
Off-platform activities: Another important source of data through which these social media platforms collect enormous amounts of digital behavioral data is off-platform activities. This is a list of all the activities in external platforms (outside the source platform) that these platforms have tracked or linked to a user. For instance, if a user visits a website or application that also installs one of the tracking techniques of any of the short-format platforms, then the said website can share the user’s activities, such as page view, purchase, search, logins, etc., along with other activities, such as adding a product to the cart, view contents, install app, launch app, etc. Instagram and TikTok DDPs contain off-platform activities of users, whereas YouTube does not have such data. Both Instagram and TikTok share the external platform name, activity type, and the timestamp.
@@
The DDPs from the three platforms also provide a variety of other details. For example, they may contain information on link history, survey, application settings, shopping, etc. For more details, please refer to Table III.
3.6 Key omissions in current implementation
While the previous sections bring out some platform specific deficiencies in the content for different categories of data, the GDPR Right of Access requires the implementation to communicate much more than just the data. For instance, Article 15(1) mandates the purpose of data collection, the third-party or categories of third-parties to whom such data has been disclosed, the period for which data will be retained, etc. Worryingly, all three platforms fail to report this information in accordance with the Right of Access. While we acknowledge that some of the information might be available in different policy-related pages of the platform, we posit that it is the obligation of the platform to make this information explicit and accessible.
Furthermore, none of the DDPs provide any README or manual to help users understand the significance of the different categories of collected data. As explained in this section, most often, platforms provide the raw data without providing any metadata or context. Without proper context, digesting and making sense of such complex information may be further difficult for common users.
Important takeaways:
Despite offering similar services, there are significant inconsistencies in data shared (and thus gathered) by the three platforms. Instagram collects disproportionately higher granularity of personal data than TikTok and YouTube.
There are inconsistencies in the reporting of data within platforms. For example, while some details (e.g., URL to the content) are available for one category of data (like history) they are missing in another category (watch history) for Instagram DDP.
All the studied platforms fail to report purpose, recipients, and retention period of collected data. These shortcomings defeat the purpose of such data protection rights.
4 Auditing reliability of DDPs
While the platforms differ in their data collection and sharing practices, as long as they reliably share the information with users it still is acceptable. Therefore, in this section, we audit the reliability of DDPs provided by the platforms along the following dimensions.
Completeness: Evaluates whether the platforms provide a comprehensive record of all user activities or the data is subject to sampling or omissions.
Correctness: Determines whether the sequence of recorded activities in the DDP accurately reflects the chronological order of events.
Consistency: Assesses the uniformity of data within same user account upon different requests and across different user accounts.
4.1 Data collection
To evaluate completeness and correctness of the data within DDPs, we need access to the ground truth i.e., what are the actual activities done on the corresponding user account. Hence, we use sock-puppet accounts to systematically browse and log the behavior of the sock-puppet to generate the ground truth. Next, we elaborate on the detailed procedure of data collection through sock-puppet accounts.
Sock-puppet accounts: We automated video browsing and liking activities using pyautogui [2] to control mouse and keyboard and opencv-python [1] to detect the like button on the desktop screen. During the activities, we captured interactions with the browser by downloading the HAR (HTTP Archive) files (HAR is a JSON-formatted file that logs all network traffic between the browser and the server). We viewed between 20 to 25 short-format videos in each video browsing session on each of the three platforms. Our bot accounts watched each video for a random duration, ranging from 15 to 60 seconds, and they randomly liked some of the videos. This approach allowed us to gather accurate ground truth data regarding user activities. With this ground-truth dataset, to assess the reliability of the platform provided DDPs, we, then, exercise the GDPR Right of Access from these accounts to get the DDPs from the platforms. This process was conducted for nearly thirty video browsing sessions between October 2024 and December 2024 for each platform to avoid any form of stochasticity in observations.
While the above data enables us to audit the correctness and completeness of DDPs, it does not fully capture the different nuances of consistency. For example, multiple time requesting for DDPs in the above approach allows us to audit consistency within an account. However, we still can not claim anything about the generalizability of the observation across accounts. Therefore, to better assess the consistency of the DDPs, we decided to collect data from real users.
Real-world users: To collect data from real users, we designed our own data donation website. To ensure user privacy, all collected data was anonymized and any PII was discarded on the front-end of our data donation website. During the data collection process, we intentionally mandated the donation of specific data categories for each platform. For TikTok and YouTube, video browsing history was mandatory. In the case of Instagram, the mandatory categories included (a) like history and (b) ads and topics, which consist of posts, ads, videos viewed, and ads clicked. Other data categories, such as search history, comments, etc., were optional and users provided them at their discretion. Private messages and other sensitive data were not collected. We recruited the users on Prolific [25] and employed the platform’s screening features to select the participants who met our eligibility criteria. Specifically, participants were required to have an approval rate of at least 98%, a minimum of 100 prior contributions and have regular engagement with Instagram, TikTok and YouTube. We also mandated that the users should have at least 90 days of activity on each of these platforms. Users were compensated \$5 for the mandatory categories and \$1 for any additional category. We collected the data from the same set of users, i.e., all the users were compensated for their data donations for all the three platforms.
Geo-locations for our data collection: First, we collected data from 10 participants from Germany, France, Italy and Spain – countries within the European Union that have the highest GDP [31]. While the primary focus of our study is to assess the implementation of Right of Access under the GDPR (which affects residents within EU), to understand the potential ‘Brussels effect’ happening in enabling Right of Access across the world, we collected data from 10 participants each from the UK, Brazil, the USA, and India [^2] For users in India, only Instagram and YouTube data were collected, as TikTok is banned in the country.. Table VIII in Appendix B shows the user demographics.
4.2 Audit methodology
@@
[…]
4.3 Observations
Completeness: As mentioned in Section 3.2, Instagram and TikTok do not provide the entire watch history of a user, since they register on the platform. Hence, these DDPs are incomplete and potentially violate Article 15(1)(d). Next, we evaluate completeness of the provided data.
Instagram offers two video browsing feeds: (1) home feed, and (2) reels feed. For Instagram’s home feed, all the activities (watching and liking) present in our recorded logs were also present in the requested DDPs. However, we observed that videos watched under the reels feed were not present in the DDP, although like activities associated with them were recorded in the DDP. This observation raises serious reliability concerns on the shared DDPs of Instagram. On TikTok, all browsing and like activities were fully recorded, with a completeness rate of 100%. For YouTube, the recorded entries accounted for 99.5% of the activities observed in the HAR dump.
Correctness: Table V presents the average Jaccard scores across the video browsing sessions for the three platforms. For Instagram, there is a drop in the Jaccard scores of browsing, because of changes in usernames, i.e, account username change or account deletion. For TikTok, Jaccard scores achieved 100% accuracy across both browsing and like activities. For YouTube, the average Jaccard scores are 100% for date, 99.83% for context, and 99.5% overall. The slight drop in the overall score is because of the 0.5% missing entries discussed earlier. For TikTok and YouTube, the recorded timestamps showed a minimal difference of $\pm$ 5 seconds compared to the HAR dump values for the same activity. However, for Instagram, the difference was as high as up to one minute, leading to a reduction in the accuracy for date and overall. Thus, all the DDPs showed higher correctness during our audit.
Platform (Activity) Date Context Overall
Instagram (Likes) 100% 100% 100%
Instagram (Browse) 96% 97% 91%
TikTok (Browse) 100% 100% 100%
TikTok (Likes) 100% 100% 100%
YouTube (Browse) 100% 99.83% 99.5%
TABLE V: Average Jaccard scores across different video browsing sessions. Concerning correctness, TikTok sits at the top, followed by YouTube and Instagram.
(a) Instagram
(b) TikTok
(c) YouTube
Consistency (intra-user):  Figure 2 presents the ratio of overlapping entries between two DDPs to the total number of entries in the earlier DDP for three key aspects, i.e., date, context, and overall. For Instagram, despite using the same accounts, we observed variations in the covered duration of the video browsing history across multiple snapshots, ranging from 6 to 13 days. Hence, due to limited availability of browsing history, we focus on analyzing the like history.  Figure 2(a) shows the results for Instagram. We observe 6% of the entries are missing from snapshot 3 to snapshot 4.
On the other hand,  Figure 2(b) illustrates that TikTok provides consistent data across multiple retrieval attempts. Lastly, in  Figure 2(c), we observe that 12% of entries were missing for snapshot pairs (1, 3) for YouTube. Among these, nearly 62% entries were advertisements, while the other videos got deleted from the platform. To sum up, while we found TikTok to be consistent in our audit, Instagram and YouTube failed our consistency audits.
Consistency (inter-user): However, these observation could be stochastic for being observed in one sock-puppet account per platform. To understand the extent to which the distribution of provided data differs across users, we analyzed the cumulative distribution function (CDF) of the durations of video browsing, search, and like history across the platforms. Figure 3 presents a comparative view of activity durations across users. For Instagram, we observe distinct clusters of users with browsing history durations concentrated around 6 and 13 days (Figure 3(a)). This non-uniform distribution has also been observed within the same user upon multiple requests over time.
On TikTok (Figure 3(b)), video browsing and search history durations exhibit substantial variation. Specifically, some end-users had data spanning approximately 180 days, while others had nearly 455 days of recorded video browsing and search activities. To ensure a fair comparison, we specifically restricted our analysis to TikTok users who have been using TikTok for at least 450 days. This criterion allows us to focus on the subset of users who have had ample time to accumulate browsing and search history. As a result, we analyzed data from 49 TikTok users. These findings highlight TikTok’s tendency to provide data inconsistently among its users. However, such clusters were not observed for like history on both Instagram and TikTok. (See Figure 10 in Appendix LABEL:appendix:reliability). The analysis of video browsing and search history on YouTube (Figure 3(c)) does not reveal any discernible patterns, suggesting that the platform may provide DDPs more consistently to its end users. This observation holds across both the EU and the other regions.
Inter-user consistency (across EU countries): Such inter-user consistency analysis shows even more interesting patterns across different countries within the EU. Instagram users in EU countries typically receive either one week or two weeks of browsing history (see Figure 4(a)). In contrast, TikTok displays considerable inconsistency: users within the same countries (Germany and France) often receive different duration of data, despite being part of the same regulatory region (see Figure 4(c)).
Inter-user consistency (across regions): We observe the ‘Brussels effect’ in terms of users in various other countries (UK, Brazil, USA and India) it is also possible to request and get a copy of personal data. Instagram continues to provide users with either one or two weeks of data across regions (see Figure 4(b)), following the same pattern observed across EU countries. On the contrary, as shown in Figure 4(d), we find that TikTok shows more uniform behavior in the USA and Brazil (around 180 days of browsing history data is provided) compared to the erratic behavior within EU.
(a) Instagram
(b) TikTok
(c) YouTube
Figure 3: Plots illustrating the CDF of data duration provided to users for two activities – browse and search history. On Instagram, clusters in browsing history are observed at 6 and 13 days, while on TikTok, clusters appear at 180 and 455 days for both search and browse history.
(a) Instagram (EU)
(b) Instagram (Regions)
(c) TikTok (EU)
(d) TikTok (Regions)
Figure 4: Plots showing the percentage of users with different browsing history durations. For Instagram, users within the same country or region were either given one week or two weeks of data. TikTok, by contrast, exhibits greater variability in the EU whereas in the USA, users were provided with the same duration of data.
Important takeaways:
TikTok DDPs are the most reliable. While, Instagram failed the completeness and consistency audits for our sock-puppet accounts, YouTube failed the consistency audit.
Instagram and YouTube exhibit varying degrees of missing data across snapshots. Moreover, Instagram’s DDP is not complete as it does not contain viewed reels.
There are disparities in the amount of data within a platform’s DDPs across users within EU. Instagram and TikTok have clear variations for video browsing and search histories, while YouTube appears to maintain a more consistent data distribution by providing data for the entire history.
5 Assessing comprehensibility of DDPs
Despite the presence of the most reliable form of information, the utility of DDPs will be significantly undermined if the communication of such information is not comprehensible to users. Therefore, in this section, we investigate whether the current implementations of platforms adhere to the GDPR’s comprehensibility requirements.
@@ 5 Assessing comprehensibility of DDPs
Methodology overview: Per Article 12 of GDPR, DDPs should be in a concise, transparent, intelligible and easily accessible form, using clear and plain language [10]. As the GDPR does not provide more succinct definitions of these requirements, following trends in empirical contract research  [4], we conduct a large-scale user survey among participants from different European countries to understand their interpretations of some of the requirements. Finally, we evaluate the current implementations, i.e., the DDPs obtained from TikTok, Instagram, and YouTube, against the interpretations by participants.
@@
Participant recruitment: To conduct our survey, we recruited 100 participants each from Germany, France, Spain and Italy (400 in total) who had a high approval rate ($\geq 98\%$) on Prolific [25]. Our motivation for a general set of respondents stems from the fact that the GDPR empowers general residents in the European Union (EU) with the Right to data access, and requirements are intended to benefit these residents. Though, eventually, our plan is to extend the study to all countries within the EU, we chose to start with the countries that have the highest GDP [31]. We also selected the standard gender breakdown while setting up the surveys. In total, we recruited 228 males (57.0%), 161 females (40.3%), 9 self-reported as “other” (2.3%), and 2 participants who preferred not to disclose their gender (0.5%) (See Table VIII in Appendix B for details). Prior to entering our survey, we presented an online consent form for expressing explicit consent from the participants. Upon completion of the survey, we compensated each participant at a rate of £9 per hour, which is recommended by Prolific to be a good and ethical rate of remuneration [24].
Our survey comprised of three components: (a) awareness of participants about the rights, (b) their interpretation of the requirements and (c) adherence evaluation. We elaborate on the survey setup and important observations for the three components.
@@ 5.2 Interpretation of the requirements
Survey setup: To understand the interpretation of requirements by common end users, we showed the participants the relevant paragraph of Article 12, which mentions the different requirements of DDPs. We asked our participants about their interpretation or expectation of the shared personal data in the DDPs to follow any of these requirements. To understand their interpretations of different requirements, we utilized BERTopic [14]. BERTopic clustered similar interpretations together into ‘topics’; we then manually derived themes based on keywords and representative sentences to understand the underlying abstract nuances. Next, we elaborate on the observations in the context of each comprehensibility requirements.
Conciseness: The important themes that resonated across the responses spanning all countries are – prioritize important & relevant data, summarize important observations, and present data in a structured way that is easier to navigate. We observe an interesting difference in the interpretations across countries. Participants from Italy, France and Spain mentioned the documents should be shorter, while others mentioned avoiding an overwhelming amount of text. In contrast, participants from Germany only prioritize the informativeness of the content irrespective of its length.
@@
Transparency: Unlike the case of conciseness, there are no major differences in how participants from all countries interpret transparency. The interpretation of transparency can be divided into two major aspects. First, participants call for unambiguous, full disclosure of all the data that has been collected about them in the way they are collected. Second, and most importantly, some participants also expect the DDPs to mention how the data is processed and for what purpose the data is collected.
Conciseness vs. Transparency: As per participants’ interpretations, conciseness and transparency requirements are at odds with each other. Therefore, any DDP representation may not satisfy these requirements simultaneously, i.e., to satisfy one, the other has to be (at least partly) disregarded.
@@
Intelligibility: All the participants explicitly interpret it to be easy to understand and link it closely with the language being straightforward and without any technical or legal jargon. Participants from all countries also mentioned that intelligible data should have a clear explanation of how and why a certain piece of information is collected. To this end, participants also noted some interesting expectations. For instance, participants from Germany mentioned that platforms should share DDPs in one’s language of choice. Moreover, participants from Italy stated that platforms should share their DDPs in the form of an elegant visualization.
@@
Clear and plain language: Most of the participants interpret plain and simple language to be words that are easily understandable and unambiguous. Some participants elaborate on the nuance by stating that legal and technical terms should be avoided or at least explained in a simpler way. To this end, we observe that participants do not distinguish between intelligibility and language requirements.
@@
Accessibility: Most of our participants interpret accessibility as being easily available, i.e., it is easy to find the place to request the data and to access the data. However, there are certain interesting and valid interpretations that are worth mentioning. Some participants mentioned that the data needs to be available in a format so that it can be opened without any sophisticated software. Also, the timeliness with which the data is provided upon request is an interesting aspect mentioned by participants. Another important interpretation was to make sure that users with visual impairments should be able to access their data using features like screen readers.
Interpretations by EDPB: To understand whether participants’ interpretations align with that of legal practitioners, we refer to the European Data Protection Board’s (EDPB) [12] guidelines which is to be followed by platforms while implementing the requirements. Along with agreeing on most of the above interpretations, EDPB also adds some more nuanced interpretations in its guidelines.
While both participants and EDPB interpret concise presentation to be succinct presentation to avoid information fatigue, EDPB also asserts platforms should differentiate between privacy sensitive and non-sensitive information in DDPs. Similarly, in addition to the mentioned interpretation of transparency, EDPB also asks platforms to clearly mention the most important consequences and risks that each data processing category may entail. Much like the tension we found between participants’ interpretation, even EDPB recommendations reflect such tensions between conciseness and transparency. For intelligibility, EDPB completely agrees with participants’ interpretations by linking it to the usage of clear and plain language which could be understood by an average member of the intended audience. Along with understandability, for the interpretation of clear and plain language, EDPB recommends using definitive language that does not leave any room for different interpretations. Finally, for accessibility, EDPB’s recommendations completely match that of our surveyd participants.